# Natural, Artificial, and Human Intelligences

### v1

@@ 6 History of (Artificial) Intelligence Tests and Moving the Goal Posts
Some of the early approaches attempted to test intelligence in terms of sensory ability (Galton,, 1883; Jensen,, 2011), an idea which was criticized by Wilhelm Wundt, who favoured complex cognitive thought, such as problem solving (Blumenthal,, 1975; Wundt,, 1896). Wundt’s ideas influenced the first standardised intelligence test, the Binet-Simon scale, and the idea of mental age, as a way to characterise developmental delays in intelligence (Binet and Simon, (1907)).
@@
The Binet-Simon scale inspired later intelligence tests, such as the Wechsler Adult Intelligence Scale (WAIS; (Wechsler,, 1939)) and, in its US reincarnation (Terman,, 1916), the Army Alpha (verbal-based) and Beta (non-verbal, for e.g. non-English speakers) tests, originally used for screening US Army recruits during the first world war (Yerkes,, 1921). This work also led to the term intelligence quotient (IQ), the ratio of mental to chronological age (Stern,, 1912). At around that time, researchers started recognising how some of these tests are culturally biased, favouring those from Western, educated, industrialised backgrounds (Bond,, 1934) . Such concerns led to the development of ostensibly culturally fair and non-verbal, such as Raven’s Progressive Matrices (Raven,, 1938) and the culture-fair intelligence test (Cattell,, 1943), an effort which continues to this date (Nisbett et al.,, 2012). An interesting assumption in these proposals is that verbal ability is largely due to education and so should not be part of intelligence testing. Today, successors to the Stanford-Binet test (Terman,, 1916) and the WAIS test are still widely used, in e.g. clinical diagnosis, education, and cognitive functioning evaluations. Pertinently, there is evidence that the more recent chatbots (such as GPT-4) can solve a variety of problems used in IQ tests, including in mathematics, verbal reasoning, comprehension of complex ideas etc. (Bubeck et al.,, 2023), as well as perform highly in standard college applicant aptitude tests, such as the Stanford Achievement Test (SAT) and the Graduate Record Examination (GRE; OpenAI, (2023)). Strong performance has already been observed with the type of problems typical in Raven’s Progressive Matrices test (Webb et al.,, 2023) and the recent versions of the WAIS. In February 2025, Grok AI announced expert-level achievements on advanced math (AIME), science (GPQA), and programming (LiveCodeBench) tests (xAI Team,, 2025). Where recent chatbots (such as GPT-4) do not perform as well, this can often be traced to technical limitations of such models (e.g., visual processing; Hernández-Orallo, (2017)).
@@
Using tests designed for human examinees to benchmark AI systems has attracted recent attention, enabled by the sophistication of today’s AI models. In previous decades, AI testing methodologies have typically focused on basic competencies, such as proportion of working transistors. Software tests have traditionally been carried out at different levels: unit testing checks that individual functions perform their calculations correctly; integration testing checks that the parts, when assembled, demonstrate the expected overall system behaviours; and regression testing checks that an updated version or specific instance of a system does not unexpectedly fall behind established performance expectations. None of these processes has been called ‘intelligence testing’, though there are analogies with the approaches of (Binet and Simon,, 1907) and educational testing since: for example, the 2 times table is such a simple test, it should be passed by all students and assumed to be known thereafter.
@@
Evaluation in machine learning has gradually adopted a more statistical approach, as outlined in Section 5. Instead of asking human experts to engineer rules for how to classify objects as a basis for (initially) relevance judgments (Cleverdon et al.,, 1966) and (later on) AI, human experts were asked to judge what the results should be. These judgements or annotations are then used as training examples for assessing AI accuracy. Examples include relevance judgments for document retrieval, what characters should be recognized in a handwriting sample, and whether an email message is spam or junk mail (Cormack et al.,, 2008). The use of standardized measures on particular datasets raises important questions, including sampling bias and whether a good benchmark result guarantees appropriate behaviour in real applications. Another challenge is that human judges do not always agree: measuring and studying inter-annotator agreement, or inter-rater reliability, is an important part of the methodology (Hallgren,, 2012), though once an annotated dataset is available, it is common for machine learning evaluations to ignore this issue, even referring to labels from human annotators as ‘ground truth’. Notwithstanding these issues, the gradual adoption of established datasets and quantitative evaluations is part of what has enabled AI to make consistent progress since the 1990s, including the use of BLEU scores to evaluate machine translation (Bahdanau et al.,, 2014; Vaswani et al.,, 2017), and the development of comprehensive sets of language processing benchmarks designed to cover a variety of competencies (Wang,, 2018).
@@
Evaluation of systems by competing directly with human performance has been less of a motivating force in AI until recently. Initially, it was obvious that computers were faster and more reliable at arithmetic calculations (this is what they were built for), so there was little point in comparing human with mechanical performance, just as we would not race a powered vehicle against a human runner. By contrast, most other human behaviours regarded as intelligent were clearly beyond early computers. Games like chess and Go became exceptions, but organized competitions for Turing-like tests (particularly the prize offering by AI inventor Hugh Loebner) have declined since the late 2010s (Singh and Beniwal,, 2022) (though just as this paper was being written, the experiments of (Jones and Bergen,, 2025) found that GPT-4.5 fooled human judges into choosing it as the ‘real’ human in an astonishing 73% of cases). The recent strong performance of AI systems on traditional intelligence tests for humans is not the culmination of decades in pursuit of this goal: it is more the result of decades of progress in lower-level benchmark evaluations, with the surprising outcome that systems trained to predict next words in a document or appropriate answers to relatively simple questions ended up performing well on tests designed for humans.
@@
(Levesque et al.,, 2012)
As emphasized by (Chollet,, 2019), a pioneer of more general tests in AI, challenges like this are designed to stimulate, not to disparage, progress in AI, and sure enough, several systems had successfully tackled the Winograd schema challenge by 2020 (Kocijan et al.,, 2023). Chollet, (2019) proposed a set of challenges called the Abstract Reasoning Corpus (ARC), which focus on shapes in grid patterns and how they might interact physically: for example, it is easy for humans to see that a red shape is moved until it collides with a blue shape and to apply the same ‘rule’ immediately to the next scene. As of early 2025, performance of AI systems on ARC tests has leaped dramatically and, while there is discussion on whether the test has been conclusively passed (due particularly to energy considerations – the official ARC test included energy-consumption limits that were breached by GPT-03), Chollet, (2024) praises the groundbreaking progress in AI.
It is easy to find evidence that connecting words with images in an exact fashion is still challenging for AI systems. State-of-the-art AI systems can generate stylish images from text prompts, but without preserving key facts: for example, a C major scale should have the notes CDEFGABC, and a chessboard at the beginning of the game should have exactly 64 squares and 32 pieces; AI-generated images still fail to meet these simple requirements (Figure 2). Such gaps are perhaps even more surprising given the achievements of these models at tasks considered hard for humans (including the skilful drawing of shape, texture, and shading). Though note, more recent chatbot versions can produce outputs already better in diagrammatic accuracy.
@@
Summarizing the apparent trends, in many IQ tests, chatbots already outperform most humans and it should not be long before this is the case for all IQ tests. GPT-4’s own estimate of when this will happen is 2030. Such demonstrations invite a criticism that AI systems can accomplish this with vastly more resources, compared to human brains. For example, GPT-4 has approximately one trillion parameters and an estimated cost of about \$100M to train, while Deep Blue had about 14,000 specialized chess chips. However, comparing complexity between human and artificial agents is not straightforward. First, the human brain is thought to have more than 80 billion neurons and 100 trillion neuronal connections (Azevedo et al.,, 2009; Herculano-Houzel,, 2012), leading some scientists to claim that it is “…the most complicated object in the known universe”. Second, the brain has optimised many hardwired structures, making it efficient in cognitive tasks related to our physical constraints, environment, and needs. This optimisation has taken place across hundreds of millions of years for some generic abilities (e.g., visual perception; the time references the emergence of the first mammals, Kemp, (2005)) and across hundreds of thousands of years for cognitive abilities more specific to humans (the time references the appearance of Homo Sapiens, Hublin et al., (2017)). So, while it might appear that some of these AI systems have frightening complexity, our brains are not exactly simple either.
@@
Another possible objection is that some of the AI solutions to particular problems may appear uncompelling, because they involve brute force approaches. For example, Deep Blue can evaluate approximately 200 million positions per second and, even though it uses various pruning algorithms to reduce calculation, it still searches through millions of moves Campbell et al., (2002). Cholet’s 2019 ARC test, intended to be the new gold standard regarding testing for AI, has recently been partially overcome, by GPT-o3 Chollet, (2024). Recent developments leading to this achievement include Chain-of-Thought prompting, which automates the process of breaking a problem into basic steps (Wei et al.,, 2022), and test-time scaling, which devotes extra computational resources at inference time to check and revise answers (Muennighoff et al.,, 2025). The techniques used by such systems could be described as successful implementations of the proverbial advice “think before you speak”. Arguably, the success of inference-time improvements is just the application of brute force computation outside the training process. However, the human brain is also an example of a brute force approach to intelligence, if not in how we solve problems, certainly in how evolution has optimised cognitive processes for the range of problems relevant to us. We suggest that the reluctance to call e.g. Deep Blue intelligent is less so about its brute force approach and more about domain specificity: we are unlikely to call intelligent an agent who excels (however exceptionally) in a single domain (cf. the distinction between skill and intelligence, (Chollet,, 2019)).
@@
Finally, it may look that, as soon as an artificial system passes the current gold standard test for intelligence, we decide that the test was inadequate and invent a more elaborate one. That is, every time an AI test is passed, we move the goal posts and change our definition of intelligence, (apparently) deliberately so as to keep out the most sophisticated AI system. A critical perspective would be that we just rig the game to retain a perception of superiority. Some such examples are noted above (e.g., chess), but there are others:having a flawless conversation with another human or translating a document into other languages would have been unquestioned signs of intelligence, until computers could do such tasks. An alternative perspective is that revisions in AI tests simply reflect an increasing understanding of what intelligence is.
@@ 7 The Relation between Embodiment and Intelligence
Human intelligence is tied to the way we perceive and manipulate objects in the world (Wilson et al.,, 2021)), as well as our cultural and social context (Cobb,, 2001). One can ask what is the value of intelligence for solving embodiment problems — we have developed a wide range of tools, which allow successful embodied problem-solving. Some researchers go as far as to suggest that all knowledge we have is either embodied or has its origins in embodied knowledge (Johnson and Lakoff,, 2008; Lakoff and Johnson,, 1999). For example, even in the most abstract scientific theories we have constructed, such as quantum mechanics, some of the basic ideas are extensions of embodied concepts (e.g., waves and particles); the most profound conceptual challenges in such theories come from ideas with no embodied analogues (e.g., superposition, non-locality; Einstein et al., (1935)).
Could one rephrase the main question in this work from “is chat GPT intelligent”, to “can chat GPT think”? However tempting the equivalence between the two questions, it is misleading, since the latter version is too anthropomorphic to be useful. This helps bring into focus the anthropocentric bias throughout the literature regarding human intelligence and AI tests. As we saw, this has been recognised, in that there has been effort in inventing ‘disembodied’ tests of intelligence, not tied to our specific embodiment constraints. However, this approach has led to an over-emphasis on intelligence tests based on language and manipulation of information through linguistic means. Chatbots expose a fatal weakness to such tests, because, even though they are supremely capable wordsmiths, we are left uncertain as to whether they are intelligent or not. Arguably, as shown in Figure 2, key problems with state-of-the-art chatbots are less so about the question “Can chatbots think?” but rather about the question “Can chatbots see what they are doing?” — that is, the problem is not about the thought process, but about basic awareness. Indeed, Figure 2 highlights major problems with state-of-the-art chatbots, which translated into human terms would not be “Can chatbots think?” but “Can multimodal AI systems see what they’re doing?”: the problem is not the quality of the thought process, but lack of perceptual ability. Note, this point is separate from recognising that some means of abstractly expressing knowledge (not necessarily human-like language) is essential for intelligence.
@@
There is no doubt that to understand human intelligence, we have to consider embodiment. Likewise, some artificial intelligence researchers have argued that artificial agents need some form of embodiment (e.g., through robotics; Anderson, (2003)), before they can develop intelligence. Regarding chatbots, given that they lack any aspect of embodiment and their only basis for learning is input from language, it is unsurprising that any intelligence they develop would be different from ours. However, ideally our notion of intelligence would be versatile enough to recognise different kinds of intelligence (e.g., based on agents embodied in different ways or no embodiment at all). That is, to understand what it means to be human and human intelligence, we need embodiment, but our notion of intelligence in general should not be tied to a particular embodiment. Of course, historically we (humans) have been appropriating the notion of general intelligence to mean human intelligence, though we (the authors) think current evidence suggests otherwise (that is, that there can be general intelligence unlike ours).
@@
The discussion of embodiment can perhaps help disentangle skill from intelligence, a challenge which arises both for human and artificial intelligence (c.f., Aristotle’s Metaphysics, Bk 1, Ch1, and (Chollet,, 2019)). We are part of a culture that has assumed, for some centuries, that penmanship is intelligence, whereas manual dexterity is “mere” skill, and the use of physical force is sometimes disparaged as brutish. A cardinal example is the phrase “the pen is mightier than the sword” — an adage celebrated even on the walls of the Library of Congress (the quote is from an 1839 play by Edward Bulwer-Lytton, in the mouth of Cardinal Richelieu, First Minister of France from 1622 to 1642). If we remove the warlike associations and propose more blandly that “the pen is more useful than the knife”, then the claim is obviously not true in general and depends on the work to be done. For example, knives did not depend on pens for their invention, but the other way round. So, is the use of a pen or a knife more indicative of human intelligence? One way to address these questions is by arguing that intelligent behaviour does not depend on whether the task involves language or mathematics or manual labour etc. Rather, the key factor is the complexity of the task.
8 Our Proposal

### v2

@@ 6 History of (Artificial) Intelligence Tests and Moving the Goal Posts
Some of the early approaches attempted to test intelligence in terms of sensory ability (galton1883inquiries; jensen2011theory), an idea which was criticized by Wilhelm Wundt, who favoured complex cognitive thought, such as problem solving (blumenthal1975reappraisal; wundt1896lectures). Wundt’s ideas influenced the first standardised intelligence test, the Binet-Simon scale, and the idea of mental age, as a way to characterise developmental delays in intelligence (binet1907enfants).
@@
The Binet-Simon scale inspired later intelligence tests, such as the Wechsler Adult Intelligence Scale (WAIS; (wechsler1939measurement)) and, in its US reincarnation (terman1916measurement), the Army Alpha (verbal-based) and Beta (non-verbal, for e.g. non-English speakers) tests, originally used for screening US Army recruits during the first world war (yerkes1921psychological). This work also led to the term intelligence quotient (IQ), the ratio of mental to chronological age (stern1912psychologischen). At around that time, researchers started recognising how some of these tests are culturally biased, favouring those from Western, educated, industrialised backgrounds (bond1934education). Such concerns led to the development of ostensibly culturally fair and non-verbal tests, such as Raven’s Progressive Matrices (raven1938progressive) and the culture-fair intelligence test (cattell1943measurement), an effort which continues to this date (nisbett2012intelligence). An interesting assumption in these proposals is that verbal ability is largely due to education and so should not be part of intelligence testing. Today, successors to the Stanford-Binet test (terman1916measurement) and the WAIS test are still widely used, in e.g. clinical diagnosis, education, and cognitive functioning evaluations. Pertinently, there is evidence that the more recent chatbots (such as GPT-4) can solve a variety of problems used in IQ tests, including in mathematics, verbal reasoning, comprehension of complex ideas etc. (bubeck2023sparks), as well as perform highly in standard college applicant aptitude tests, such as the Stanford Achievement Test (SAT) and the Graduate Record Examination (GRE; openai2023gpt). Strong performance has already been observed with the type of problems typical in Raven’s Progressive Matrices test (webb2023emergent) and the recent versions of the WAIS. In February 2025, Grok AI announced expert-level achievements on advanced math (AIME), science (GPQA), and programming (LiveCodeBench) tests (xai2025grok3beta). Where recent chatbots (such as GPT-4) do not perform as well, this can often be traced to technical limitations of such models (e.g., visual processing; hernandez-orallo2017measure).
@@
Using tests designed for human examinees to benchmark AI systems has attracted recent attention, enabled by the sophistication of today’s AI models. In previous decades, AI testing methodologies have typically focused on basic competencies, such as proportion of working transistors. Software tests have traditionally been carried out at different levels: unit testing checks that individual functions perform their calculations correctly; integration testing checks that the parts, when assembled, demonstrate the expected overall system behaviours; and regression testing checks that an updated version or specific instance of a system does not unexpectedly fall behind established performance expectations. None of these processes has been called ‘intelligence testing’, though there are analogies with the approaches of (binet1907enfants) and educational testing since: for example, the 2 times table is such a simple test, it should be passed by all students and assumed to be known thereafter.
@@
Evaluation in machine learning has gradually adopted a more statistical approach, as outlined in Section 5. Instead of asking human experts to engineer rules for how to classify objects as a basis for (initially) relevance judgments (cleverdon1966factors) and (later on) AI, human experts were asked to judge what the results should be. These judgements or annotations are then used as training examples for assessing AI accuracy. Examples include relevance judgments for document retrieval, what characters should be recognized in a handwriting sample, and whether an email message is spam or junk mail (cormack2008email). The use of standardized measures on particular datasets raises important questions, including sampling bias and whether a good benchmark result guarantees appropriate behaviour in real applications. Another challenge is that human judges do not always agree: measuring and studying inter-annotator agreement, or inter-rater reliability, is an important part of the methodology (hallgren2012computing), though once an annotated dataset is available, it is common for machine learning evaluations to ignore this issue, even referring to labels from human annotators as ‘ground truth’. Notwithstanding these issues, the gradual adoption of established datasets and quantitative evaluations is part of what has enabled AI to make consistent progress since the 1990s, including the use of BLEU scores to evaluate machine translation (bahdanau2014neural; vaswani2017attention), and the development of comprehensive sets of language processing benchmarks designed to cover a variety of competencies (wang2018glue).
@@
Evaluation of systems by competing directly with human performance has been less of a motivating force in AI until recently. Initially, it was obvious that computers were faster and more reliable at arithmetic calculations (this is what they were built for), so there was little point in comparing human with mechanical performance, just as we would not race a powered vehicle against a human runner. By contrast, most other human behaviours regarded as intelligent were clearly beyond early computers. Games like chess and Go became exceptions, but organized competitions for Turing-like tests (particularly the prize offering by AI inventor Hugh Loebner) have declined since the late 2010s (singh2022survey) (though just as this paper was being written, the experiments of (jones2025large) found that GPT-4.5 fooled human judges into choosing it as the ‘real’ human in an astonishing 73% of cases). The recent strong performance of AI systems on traditional intelligence tests for humans is not the culmination of decades in pursuit of this goal: it is more the result of decades of progress in lower-level benchmark evaluations, with the surprising outcome that systems trained to predict next words in a document or appropriate answers to relatively simple questions ended up performing well on tests designed for humans.
@@
As emphasized by (chollet2019measure), a pioneer of more general tests in AI, challenges like this are designed to stimulate, not to disparage, progress in AI, and sure enough, several systems had successfully tackled the Winograd schema challenge by 2020 (kocijan2023defeat). chollet2019measure proposed a set of challenges called the Abstract Reasoning Corpus (ARC), which focus on shapes in grid patterns and how they might interact physically: for example, it is easy for humans to see that a red shape is moved until it collides with a blue shape and to apply the same ‘rule’ immediately to the next scene. As of early 2025, performance of AI systems on ARC tests has leaped dramatically and, while there is discussion on whether the test has been conclusively passed (due particularly to energy considerations – the official ARC test included energy-consumption limits that were breached by GPT-03), chollet2024o3 praises the groundbreaking progress in AI.
It is easy to find evidence that connecting words with images in an exact fashion is still challenging for AI systems. State-of-the-art AI systems can generate stylish images from text prompts, but without preserving key facts: for example, a C major scale should have the notes CDEFGABC, and a chessboard at the beginning of the game should have exactly 64 squares and 32 pieces; AI-generated images still fail to meet these simple requirements, even though they can be very creative (Figure 2). Such gaps are perhaps even more surprising given the achievements of these models at tasks considered hard for humans (including the skilful drawing of shape, texture, and shading). Though note, more recent chatbot versions can produce outputs already better in diagrammatic accuracy.
@@
Summarizing the apparent trends, in many IQ or aptitude tests, chatbots already outperform most humans. Such demonstrations invite a criticism that AI systems can accomplish this with vastly more resources, compared to human brains. For example, GPT-4 has approximately one trillion parameters and an estimated cost of about \$100M to train, while Deep Blue had about 14,000 specialized chess chips. However, comparing complexity between human and artificial agents is not straightforward. First, the human brain is thought to have more than 80 billion neurons and 100 trillion neuronal connections (azevedo2009equal; herculano-houzel2012remarkable), leading some scientists to claim that it is “…the most complicated object in the known universe”. Second, the brain has optimised many hardwired structures, making it efficient in cognitive tasks related to our physical constraints, environment, and needs. This optimisation has taken place across hundreds of millions of years for some generic abilities (e.g., visual perception; the time references the emergence of the first mammals, kemp2005origin) and across hundreds of thousands of years for cognitive abilities more specific to humans (the time references the appearance of Homo Sapiens, hublin2017new). So, while it might appear that some of these AI systems have frightening complexity, our brains are not exactly simple either, and use vastly more neurons that would be logically necessary to solve some problems.
@@
Another possible objection is that some of the AI solutions to particular problems may appear uncompelling, because they involve brute force approaches. For example, Deep Blue can evaluate approximately 200 million positions per second and, even though it uses various pruning algorithms to reduce calculation, it still searches through millions of moves campbell2002deep. Cholet’s chollet2019measure ARC test, intended to be the new gold standard regarding testing for AI, has recently been partially overcome, by GPT-o3 chollet2024o3. Recent developments leading to this achievement include Chain-of-Thought prompting, which automates the process of breaking a problem into basic steps (wei2022chain), and test-time scaling, which devotes extra computational resources at inference time to check and revise answers (muennighoff2025s1). The techniques used by such systems could be described as successful implementations of the proverbial advice “think before you speak”. Arguably, the success of inference-time improvements is just the application of brute force computation outside the training process. However, the human brain is also an example of a brute force approach to intelligence, if not in how we solve problems, certainly in how evolution has optimised cognitive processes for the range of problems relevant to us.
We suggest that the reluctance to call e.g. Deep Blue intelligent is less so about its brute force approach and more about domain specificity: we are unlikely to call intelligent an agent who excels (however exceptionally) in a single domain (cf. the distinction between skill and intelligence (chollet2019measure)). We do not call a calculator intelligent, but a small child who could multiply 4-digit numbers in their head would be considered an intellectual prodigy. A teenager who wins a spelling-bee competition is often described as ‘smart’, and while any literate human could do just as well by using a dictionary, we expect that the spelling-bee champion is also good at remembering other things. Estimating (and sometimes overestimating) the extent to which success on one test demonstrates aptitude for other activities has long been recognized as a challenge for test-driven assessments (gardner2011frames).
@@
It may look that, as soon as an artificial system passes the current gold standard test for intelligence, we decide that the test was inadequate and invent a more elaborate one. That is, every time an AI test is passed, we move the goal posts and change our definition of intelligence, (apparently) deliberately so as to keep out the most sophisticated AI system. A critical perspective would be that we just rig the game to retain a perception of superiority. Some such examples are noted above (e.g., chess), but there are others: having a flawless conversation with another human or translating a document into other languages would have been unquestioned signs of intelligence, until computers could do such tasks. An alternative perspective is that revisions in AI tests simply reflect an increasing understanding of what intelligence is.
Our belief that machines are likely to outpace humans in all clearly-definable tasks is partly because of human ingenuity and partly due to our lack of scruples (at least so far) when dealing with non-humans. Pulling open neural networks and language models to find how best to improve them is always an option, which gives us options for improvement that we can explore systematically. (This advantage is alluded to already in lecun1989optimal with the paper title “Optimal Brain Damage” — something we can explore in neural nets and not people.)
A possible limitation in considering chatbot capabilities in intelligence testing is a historic focus on the English language. For example, early language models tended to use English, as it has been in English that most data has been readily available. This has led to claims that AI agents are poor at interlingual communication. For example, “AIs are usually programmed only in a single language” (nikolarea2024can). However, when such problems are recognized within the AI field, they often become superseded. With appropriate training data, small multilingual embedding spaces have been built since the work of widdows2002parallel. Since the launch of BERT and mBERT (devlin2018bert), multilingual spaces have been standard (pires2019multilingual). Returning to the point about ’dissection’, nowadays, we can look inside many different embedding spaces in different layers, enabling machines to cooperate with one another by exchanging non-linguistic tokens from latent spaces (dessi2021interpretable); one of the most pressing research challenges is to make such automatically-derived interlingua meaningful to humans (karten2023interpretable). With AI models, we can explore, change, and experiment with every parameter in every layer, and, while modern techniques such as fMRI scans have given some insight on activity in the human brain, there is no comparison with how transparently we can see and change each individual activity in AI models.
The ability to monitor different levels of activity in AI models also bears on questions of self-awareness. The notion that an AI system cannot be self-aware like a human is one of the objections considered by turing1950computing, who regards it as in fact an appeal to solipsism (self-awareness is a personal experience, hence not something an individual can claim on behalf of a species). There is some work arguing that modern chatbots have an impressive range of metacognitive abilities (li2025aiawareness; but see also nikolarea2022translators).
However, more generally, instead of asking whether chatbots can experience self-awareness, we can focus on their behaviours. When asked about a conversation, chatbots are increasingly able to explain why they responded in a particular way, and can sometimes recommend improvements in their own configuration to avoid problems recurring. (For example, they can suggest updates to AGENTS.md files, which are used by developers to control coding agents.) From an engineering point-of-view, it does not matter whether the agent experiences self-awareness: but if it can monitor and improve its own behaviour, it becomes self-aware in the way that matters to its collaborators. As with other tasks and evaluations, if self-awareness is measured by answering introspective questions, AI will develop a foreseeable advantage, partly because we have more options for engineering the correct behaviour with chatbots than we do with ourselves.
@@ 7 The Relation between Embodiment and Intelligence
Human intelligence is tied to the way we perceive and manipulate objects in the world (wilson2021embodied)), as well as our cultural and social context (cobb2001situated). One can ask what is the value of intelligence for solving embodiment problems — we have developed a wide range of tools, which allow successful embodied problem-solving. Some researchers go as far as to suggest that all knowledge we have is either embodied or has its origins in embodied knowledge (johnson2008metaphors; lakoff1999philosophy). For example, even in the most abstract scientific theories we have constructed, such as quantum mechanics, some of the basic ideas are extensions of embodied concepts (e.g., waves and particles); the most profound conceptual challenges in such theories come from ideas with no embodied analogues (e.g., superposition, non-locality; einstein1935can).
Physical embodiment is certainly key to human intelligence (that is, our particular embodiment is tied to our notion of intelligence) and many aspects of language reflect metaphorical extension of meaning from embodied concepts. For example, as soon as we say a problem is hard and takes a long time to solve, we demonstrate a spatial grounding to our concepts (torresmartinez2022metaphors; widdows2004geometry, Ch 1). Nonetheless, chatbots are adept at using these words: it is a great surprise that artificial agents talk as if they understand the world, even when their experience is reduced to digitized text (summerfield2024strange).
Could one rephrase the main question in this work from “Are chatbots intelligent”, to “can chatbots think”? Turing predicted that “at the end of the century the use of words and general educated opinion will have altered so much that one will be able to speak of machines thinking without expecting to be contradicted.” turing1950computing Turing’s timeline estimate turns out to have been optimistic, but a quarter of the way through the next century, we do have models that use various “thinking modes”, and we are used to messages saying “Thought for X seconds”. The processes themselves are partly inspired by human approaches to problem solving, for example, by breaking reasoning down into steps and checking each step wei2022chain. Following technology history, we can expect that a chatbot thinking something will be as commonplace as a book saying something. The objection that it is not really thinking will remain active, and this concern remains important: the situation is similar to Plato’s concerns in the Phaedrus dialogue, which urge the readers themselves to remember that reading something is no substitute for truly knowing something.
As Gardner’s account of musical intelligence reminds us, not all thinking takes linguistic form (gardner2011frames). In Phaedrus, Plato joins λογος with rhythm, memory, and harmony as qualities of a well-ordered soul; learning “by heart” and reasoning through sound or motion are ways of knowing, and Plato places them closer to the psyche (ψυχη, often translated as soul) than the written word. These attributes vitally involve embodiment: so the importance of embodiment is deeply established in Western philosophy, but has become de-emphasized.
This helps bring into focus the anthropocentric bias throughout the literature regarding human intelligence and AI tests. As we saw, this has been recognised, in that there has been effort in inventing ‘disembodied’ tests of intelligence, not tied to our specific embodiment constraints. (We have also shown that this is not a universal approach, but one that has been dominant in industrialized Western societies since the 1800s.) This approach has led to an over-emphasis on intelligence tests based on language and manipulation of information through linguistic means. Chatbots expose a fatal weakness to such tests, because, even though they are supremely capable wordsmiths, we are left uncertain as to whether they are intelligent or not and wondering if humans should be considered as genuinely inferior if chatbots beat us at reasoning challenges. This tension reflects what torresmartinez2025dehumanizing calls the dialectic of dehumanizing the human while humanizing the machine, a pattern that complicates how consciousness and agency are attributed across biological and artificial systems.
Arguably, as shown in Figure 2, key problems with state-of-the-art chatbots are less so about the question “Can chatbots think?” but rather about the question “Can chatbots see what they are doing?” — that is, the problem is not about the thought process, but about basic awareness. Indeed, Figure 2 highlights major problems with state-of-the-art chatbots, which translated into human terms would not be “Can chatbots think?” but “Can multimodal AI systems see what they’re doing?”: the problem is not the quality of the thought process, but lack of perceptual ability.
This challenge has been well-recognized and addressed during 2025, by vision-aware models including Microsoft Omniparser lu2024omniparser: it is already the case that state-of-the art chatbot tools for coding include interfaces for sharing images including screenshots. Limitations on the chatbots’ contextual awareness remain, but are partly by-design, for example, the chatbot might be restricted from seeing menu options that elevate its own permissions. Some vision capabilities are still in their infancy, but we are getting used to AI agents ‘seeing’. (We are already used to voice tools like Siri and Alexa ‘hearing’.)
@@
It follows that to understand human intelligence, we have to consider embodiment. Likewise, some artificial intelligence researchers have argued that artificial agents need some form of embodiment (e.g., through robotics; anderson2003embodied), before they can develop intelligence. Regarding chatbots, given that they lack any aspect of embodiment and their only basis for learning is input from language, it is unsurprising that any intelligence they develop would be different from ours. However, ideally our notion of intelligence would be versatile enough to recognise different kinds of intelligence (e.g., based on agents embodied in different ways or no embodiment at all). That is, to understand what it means to be human and human intelligence, we need embodiment, but our notion of intelligence in general should not be tied to a particular embodiment. Of course, historically we (humans) have been appropriating the notion of general intelligence to mean human intelligence, though we (the authors) think current evidence suggests otherwise (that is, that there can be general intelligence unlike ours).
@@
The discussion of embodiment can perhaps help disentangle skill from intelligence, a challenge which arises both for human and artificial intelligence (c.f., Aristotle’s Metaphysics, Bk 1, Ch1, and (chollet2019measure)). We are part of a culture that has assumed, for some centuries, that wordsmithing is intelligence, whereas manual dexterity is “mere” skill, and the use of physical force is sometimes disparaged as brutish. A cardinal example is the phrase “the pen is mightier than the sword” — an adage celebrated even on the walls of the Library of Congress (the quote is from an 1839 play by Edward Bulwer-Lytton, in the mouth of Cardinal Richelieu, First Minister of France from 1622 to 1642). If we remove the warlike associations and propose more blandly that “the pen is more useful than the knife”, then the claim is obviously not true in general and depends on the work to be done. For example, knives did not depend on pens for their invention, but the other way round.
So, is the use of a pen or a knife more indicative of human intelligence? During the period since the 1800s, when ‘intelligence’ has been defined and measured by success at different tasks, those tasks have tended to measure skills related to Gardner’s linguistic and logico-mathematical intelligences (gardner2006multiple), often ignoring skills related to (say) bodily-kinesthetic or personal intelligences. One way to address these questions is by arguing that intelligent behaviour does not depend on whether the task involves language or mathematics or manual labour etc. Rather, the key factor is the complexity of the task. And, as noted by kasparov2017deep, we have a long history of overestimating the complexity of intellectual tasks (such as remembering chess positions), based on the difficulty most of us experience with this, compared with the ease with which we can make a cup of tea.
8 Concluding Comments